OpenStudy (anonymous):

help with problems medal given. have the answers just solve the problem 1. square root x-8=2 2. square root s+13= square root 29-s 3. x= square root 2x +24 5. y= square root 7x-4

OpenStudy (anonymous):

in num 1 just squared both sides to remove square root (square root x-8)^2=2^2 x-8=4 x=4+8

OpenStudy (anonymous):

just the same way in all num 2 square root s+13= square root 29-s squared both sides then if we squared square roots it will cancel so s+13=29-s s+s=29-13 2s=16 divide by 2 both sides s=

OpenStudy (anonymous):

\[x=\sqrt{2x+24}\] square and get \[x^2=2x+24\] or \[x^2-2x-24=0\] solve by factoring then check the result

OpenStudy (anonymous):

you get \[(x+4)(x-6)=0\] making \(x=-4\) or \(x=6\) substitute back in the original equation, you will see that \(6\) works because \[6=\sqrt{2\times 6+24}\] is true, but \(-4\) does not work

OpenStudy (anonymous):

thankk you so much @satellite73 you really helped with these simple questions which other ppl couldnt do!